OpenStudy (anonymous):

any one knows whats 11x121=

OpenStudy (anonymous):

1331?

OpenStudy (diyadiya):

1331

OpenStudy (anonymous):

R u supposed to do this by hand?

OpenStudy (anonymous):

U asked the same question twice, why?

OpenStudy (anonymous):

what do u want to lern especially?

OpenStudy (anonymous):

11 multiples of 121 = 11*121 = 11(121) = 121 + 1210 =1331

OpenStudy (anonymous):

This question has already been answered a couple posts back.

OpenStudy (anonymous):

cool thing is it looks like binomial expansion

OpenStudy (anonymous):

\[11^0=1\] \[11^1=11\] \[11^2=121\] \[11^3=1331\] \[11^4=14641\]

OpenStudy (anonymous):

another cool thing is how easy it is to multiply a number by 11 just copy it down twice, with the second one one place to the left 11 * 1243 1 2 5 3 1 2 5 3 ______________________ 1 3 7 8 3

OpenStudy (anonymous):

isn't the answer the following? 1243 1243 ------ 13,673